OpenStudy (anonymous):

Let f(x, y) = integral(1, x) of P(t) dt + integral(1, y) Q(t) dt. Find df/dx, df/dy

OpenStudy (anonymous):

So \[f(x,y) = \int\limits\limits_{1}^{x} P(t) dt +\int\limits\limits_{1}^{y} Q(t) dt\] Use the First Fundamental Theorem of Calculus: if \[g(x)=\int\limits_{a}^{x}h(t)dt\] then \[\frac{d \ g(x)}{d \ x}=h(x)\] So in this particular problem we get: \[\frac{df}{dx}=P(x)+0\] the zero is because the second integral is a function of y and dy/dx is 0. \[\frac{df}{dy}=0+Q(y)\]